OpenStudy (anonymous):

A triangle has side lengths 20, 21, and 40. Is the triangle acute, right or obtuse? Answer a. Acute b. Obtuse c. Right d. Not a triangle

OpenStudy (anonymous):

Can someone explain how to do that?

OpenStudy (cwrw238):

it is a triangle because any 2 lines is longer than the other line 20 + 21 = 41 - just 1 longer than 40 - so it will be obtuse

OpenStudy (anonymous):

ok thanks now i understand

OpenStudy (cwrw238):

looks something like 

OpenStudy (cwrw238):

yw

hero (hero):

Hint: a^2 + b^2 > c^2 = acute a^2 + b^2 < c^2 = obtuse

hero (hero):

You can't use @cwrw238's method because any two sides of a triangle is always longer than the third side no matter what

OpenStudy (cwrw238):

your equations are of course correct hero but this as an extreme case so my method did work in this scenario

hero (hero):

How is this an extreme case? It's just a triangle like any other. It will either be acute or obtuse.

hero (hero):

Your method is indeed flawed and not the logic that is used to solve these

OpenStudy (cwrw238):

hmmm

hero (hero):

Okay for example, you have sides 6,7,8

OpenStudy (anonymous):

could you help with me latest problem?

hero (hero):

According to your method, since 6 + 7 = 13 and 13 > 8 , then that means the triangle is obtuse right? @cwrw238

hero (hero):

But it is in fact acute

OpenStudy (cwrw238):

no - the 21 and 20 sides have a sum which is only 1 more than the 40 and they are only about half the length of 40 so it is obvious that the angle as drawn will be obtuse!! thats what i meant when i said it was an extreme case.

hero (hero):

Either way, your method is inconsistent and should not be used to determine these. I'm pretty sure I can find a counter example to your claim

OpenStudy (cwrw238):

yours is the correct general method but in this case I just used plain common sense - its as simple as that